OpenStudy (anonymous):

Limit as x approaches 9. (Sqrt(x)-3)/(x-9)

OpenStudy (jhannybean):

\[\lim_{x \rightarrow 9} \frac{\sqrt{x-3}}{x-9}\] is this it?

OpenStudy (jhannybean):

Or is it \[\lim_{x \rightarrow 9} \frac{\sqrt{x}-3}{x-9}\]

OpenStudy (anonymous):

its the second one

OpenStudy (anonymous):

sorry i was on my phone and it didnt let me reply and i couldnt remember my password

OpenStudy (jhannybean):

What is another way you can write out `x-9`?

OpenStudy (jhannybean):

Remember, this is a difference of squares.

OpenStudy (jhannybean):

\[x = \sqrt{x} \cdot \sqrt{x}\]\[9 = 3 \cdot 3\]

OpenStudy (jhannybean):

And difference of squares, \((a-b)^2 = (a-b)(a+b)\)

OpenStudy (anonymous):

so it's \[(\sqrt{x}-3)^2\]

OpenStudy (anonymous):

?

OpenStudy (jhannybean):

Yep :)

OpenStudy (anonymous):

then id cancel out one and id be left with 1/sqrt(9)+3 making it 6 so it'd be 1/6

OpenStudy (anonymous):

awesome thank you :)

OpenStudy (jhannybean):

So you would have : \[\lim_{x \rightarrow 9} \frac{\sqrt{x}-3}{(\sqrt{x}-3)^2}\]

OpenStudy (jhannybean):

This can be simplified how?

OpenStudy (anonymous):

by separating the bottom?

OpenStudy (jhannybean):

Mmhmm, and what does that factor out to according to the difference of squares?

OpenStudy (anonymous):

\[(\sqrt{x}-3)(\sqrt{x}-3) \]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

wait no a plus in one of them?

OpenStudy (jhannybean):

Yep, onehas a plus.

OpenStudy (jhannybean):

So therefore you would have \[\lim_{x \rightarrow 9} \frac{\color{red}{\sqrt{x}-3}}{(\color{red}{\sqrt{x}-3})(\sqrt{x}+3)}\]

OpenStudy (jhannybean):

Do you see how the ones highlighted in red can be reduced to 1?

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{9}-3 }\]

OpenStudy (anonymous):

yes, then it'd be that

OpenStudy (anonymous):

\[\frac{ 1 }{ 6 }\]

OpenStudy (jhannybean):

NOt quite, theres an error in your sign

OpenStudy (anonymous):

whats the error?

OpenStudy (anonymous):

OH the plus sign

OpenStudy (jhannybean):

\(\color{blue}{\text{Originally Posted by}}\) @Calcisnotforme2_ \[\frac{ 1 }{ \sqrt{9}\color{red}{-}3 }\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (campbell_st):

just a quick check \[(a - b)^2 \neq(a -b)(a+b)\] as stated earlier

OpenStudy (anonymous):

it should be a plus sign*

OpenStudy (jhannybean):

Oh, true campbell, thanks for identifying that. Rather, it is \(a^2-b^2 = (a-b)(a+b)\)

OpenStudy (jhannybean):

And yes @Calcisnotforme2_ it's a plus sign

OpenStudy (anonymous):

im just confused now after the note from campbell

OpenStudy (jhannybean):

Correction from earlier: \((a-b)^2 = a^2-2ab+b^2\), sorry if that confused you.

OpenStudy (anonymous):

But x-9 isn't squared. that is what im confused about. to get it to factor the way we factored it.

OpenStudy (jhannybean):

Let me clarify the two identities. and then i'll clarify which one i used to simplify the denominator.

OpenStudy (jhannybean):

1. \((a-b)^2 = \sf \text{difference of squares} = a^2 -2ab+b^2\) 2. \(\sf a^2 - b^2 = (a-b)(a+b)\)

OpenStudy (jhannybean):

If you expand the second one, you will end up with \(\sf a^2 -ab +ab - b^2 \iff a^2 -b^2\)

OpenStudy (jhannybean):

Therefore, we use the second identity to solve for \(\sf x-9\) since it can be rewritten as \(\sf (\sqrt{x})^2 -(3)^2\).

OpenStudy (jhannybean):

Does that clarify things a bit more?

OpenStudy (anonymous):

Oh okay i get it now, thank you :)

OpenStudy (jhannybean):

Yay :) You're welcome :D